One leg of a right triangle is 8 m long. The hypotenuse is 14 m long. What is the area of the triangle? Round to the nearest tenth.
Answer by checkley77(12844)

a^2+b^2=c^2
8^2+b^2=14^2
64+b^2=196
b^2=196-64
b^2=132
b=sqrt132
b=11.5 ans. for the other leg.
A=BH/2
A=8*11.5/2
A=92/2=46 m^2 is the area.